# Diffusion Bridge Implicit Models

### v1

@@ 5.2 Reconstruction and Interpolation
As discussed in Section 4.2, the deterministic nature of DBIMs at $\eta=0$ and its connection to neural ODEs enable faithful encoding and reconstruction by treating the booting noise as the latent variable. Furthermore, employing spherical linear interpolation in the latent space and subsequently decoding back to the image space allows for semantic image interpolation in image translation and image restoration tasks. These capabilities cannot be achieved by DBIMs with $\eta>0$, or by DDBM’s hybrid sampler which incorporates stochastic steps. We showcase the encoding and decoding results in Figure 5, indicating that accurate reconstruction is achievable with a sufficient number of sampling steps. We also illustrate the interpolation process in Figure 5.
@@ 6 Conclusion
In this work, we introduce diffusion bridge implicit models (DBIMs) for accelerated sampling of DDBMs without extra training. In contrast to DDBM’s continuous-time generation processes, we concentrate on discretized sampling steps and propose a series of novel generative probabilistic models on them, spanning from Markovian to non-Markovian variants. The induced sampling procedures serve as bridge counterparts and extensions of DDIMs, filling the missing perspectives of diffusion bridges. Experiments on high-resolution datasets and challenging inpainting tasks demonstrate DBIM’s superiority in both the sample quality and sample efficiency, achieving state-of-the-art FID scores with 100 steps and providing up to 25$\times$ acceleration of DDBM’s sampling procedure.
Limitations and broader impact Despite the notable speed-up for diffusion bridges, DIBMs still lag behind GAN-based methods in one-step generation, and are not fast enough for real-time applications. The efficiency of DBIMs can reduce computational resources and energy consumption in deploying diffusion bridge models. However, they may also facilitate the potential misuse for generating fake and malicious media content, and amplify undesirable social bias that could already exist in the training dataset.
Acknowledgements
@@ Appendix A Related Work
Fast sampling of diffusion models can be classified into training-free and training-based methods. A prevalent training-free fast sampler is the denoising diffusion implicit models (DDIMs) [40] that employ alternative non-Markovian generation processes in place of DDPMs, a discrete-time diffusion model. ScoreSDE [44] further links discrete-time DDPMs to continuous-time score-based models, unrevealing the generation process to be ordinary and stochastic differential equations (ODEs and SDEs). Subsequent training-free samplers concentrate on developing dedicated numerical solvers to the diffusion ODE or SDE, particularly Heun’s methods [21] and exponential integrators [49, 28, 29, 51, 12]. These methods typically require around 10 steps for high-quality generation. In contrast, training-based methods, particularly adversarial distillation [36] and consistency distillation [42, 23], become notable for their ability to achieve high-quality generation with just one or two steps. Developing dedicated differential equation solvers for DDBMs, or exploring bridge distillation methods, would be promising avenues for future research.
@@
Diffusion bridges [8, 25, 39, 53, 6, 37] are a promising generative variant of diffusion models for modeling the transport between two arbitrary distributions. One line of work is the diffusion Schrodinger bridge models [8, 37], which solves an entropy-regularized optimal transport problem between two probability distributions. However, their reliance on expensive iterative procedures has limited their application scope, particularly for high-dimensional data. Subsequent works have endeavored to enhance the tractability of the Schrodinger bridge problem by making assumptions such as paired data [25, 39, 6]. On the other hand, DDBMs [53] construct diffusion bridges via Doob’s $h$-transform, offering a reverse-time perspective of a diffusion process conditioned on given endpoints. This approach aligns the design spaces and training algorithms of DDBMs closely with those of score-based generative models, leading to state-of-the-art performance in image translation tasks. However, the sampling procedure of DDBMs still relies on inefficient simulations of differential equations, lacking theoretical insights to develop efficient samplers.
@@ Appendix B Proofs
From [3] (2.115), $q^{(\rho)}(\bm{x}_{t_{k-1}}|\bm{x}_{0},\bm{x}_{T})$ is a Gaussian, denoted as $\mathcal{N}(\bm{\mu}_{k-1},\Sigma_{k-1})$, where
@@
Substituting Eqn. (10) and Eqn. (14) into Eqn. (15), we have
@@
Appendix C Markov Property of the Generalized Diffusion Bridges
We aim to analyze the Markov property of the forward process corresponding to our generalized diffusion bridge in Section 3.1. The forward process is expressed as
@@ Appendix C Markov Property of the Generalized Diffusion Bridges
where $q^{(\rho)}(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{t_{n+1}},\bm{x}_{T})$ is defined in Eqn. (11) as
@@
Appendix D Experiment Details
D.1 Model Details

### v2

@@ 5.2 Reconstruction and Interpolation
(a) Encoding/Reconstruction
(b) Semantic Interpolation
Figure 4: Illustration of generation diversity with deterministic DBIMs.
As discussed in Section 4.2, the deterministic nature of DBIMs at $\eta=0$ and its connection to neural ODEs enable faithful encoding and reconstruction by treating the booting noise as the latent variable. Furthermore, employing spherical linear interpolation in the latent space and subsequently decoding back to the image space allows for semantic image interpolation in image translation and image restoration tasks. These capabilities cannot be achieved by DBIMs with $\eta>0$, or by DDBM’s hybrid sampler which incorporates stochastic steps. We showcase the encoding and decoding results in Figure 4a, indicating that accurate reconstruction is achievable with a sufficient number of sampling steps. We also illustrate the interpolation process in Figure 4b.
@@ 6 Conclusion
In this work, we introduce diffusion bridge implicit models (DBIMs) for accelerated sampling of DDBMs without extra training. In contrast to DDBM’s continuous-time generation processes, we concentrate on discretized sampling steps and propose a series of generalized diffusion bridge models including non-Markovian variants. The induced sampling procedures serve as bridge counterparts and extensions of DDIMs and are further extended to develop high-order numerical solvers, filling the missing perspectives in the context of diffusion bridges. Experiments on high-resolution datasets and challenging inpainting tasks demonstrate DBIM’s superiority in both the sample quality and sample efficiency, achieving state-of-the-art FID scores with 100 steps and providing up to 25$\times$ acceleration of DDBM’s sampling procedure.
Acknowledgments
@@ Appendix A Related Work
Fast sampling of diffusion models can be classified into training-free and training-based methods. A prevalent training-free fast sampler is the denoising diffusion implicit models (DDIMs) (Song et al., 2021a) that employ alternative non-Markovian generation processes in place of DDPMs, a discrete-time diffusion model. ScoreSDE (Song et al., 2021c) further links discrete-time DDPMs to continuous-time score-based models, unrevealing the generation process to be ordinary and stochastic differential equations (ODEs and SDEs). Subsequent training-free samplers concentrate on developing dedicated numerical solvers to the diffusion ODE or SDE, particularly Heun’s methods (Karras et al., 2022) and exponential integrators (Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a; Gonzalez et al., 2023). These methods typically require around 10 steps for high-quality generation. In contrast, training-based methods, particularly adversarial distillation (Sauer et al., 2023) and consistency distillation (Song et al., 2023b; Kim et al., 2023), become notable for their ability to achieve high-quality generation with just one or two steps. Developing dedicated differential equation solvers for DDBMs, or exploring bridge distillation methods, would be promising avenues for future research.
@@
Diffusion bridges (De Bortoli et al., 2021; Liu et al., 2023; Somnath et al., 2023; Zhou et al., 2023; Chen et al., 2023; Shi et al., 2024) are a promising generative variant of diffusion models for modeling the transport between two arbitrary distributions. One line of work is the diffusion Schrodinger bridge models (De Bortoli et al., 2021; Shi et al., 2024), which solves an entropy-regularized optimal transport problem between two probability distributions. However, their reliance on expensive iterative procedures has limited their application scope, particularly for high-dimensional data. Subsequent works have endeavored to enhance the tractability of the Schrodinger bridge problem by making assumptions such as paired data (Liu et al., 2023; Somnath et al., 2023; Chen et al., 2023). On the other hand, DDBMs (Zhou et al., 2023) construct diffusion bridges via Doob’s $h$-transform, offering a reverse-time perspective of a diffusion process conditioned on given endpoints. This approach aligns the design spaces and training algorithms of DDBMs closely with those of score-based generative models, leading to state-of-the-art performance in image translation tasks. However, the sampling procedure of DDBMs still relies on inefficient simulations of differential equations, lacking theoretical insights to develop efficient samplers.
@@ Appendix B Proofs
From (Bishop & Nasrabadi, 2006) (2.115), $q^{(\rho)}(\bm{x}_{t_{k-1}}|\bm{x}_{0},\bm{x}_{T})$ is a Gaussian, denoted as $\mathcal{N}(\bm{\mu}_{k-1},\Sigma_{k-1})$, where
@@
Substituting Eqn. (10) and the joint distribution into Eqn. (13), we have
@@
Appendix C More Theoretical Discussions
C.1 Markov Property of the Generalized Diffusion Bridges
We aim to analyze the Markov property of the forward process corresponding to our generalized diffusion bridge in Section 3.1. The forward process of $q^{(\rho)}$ can be induced by Bayes’ rule as
@@ C.1 Markov Property of the Generalized Diffusion Bridges
where $q^{(\rho)}(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})=q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})$ is the marginal distribution of the diffusion bridge in Eqn. (6), and $q^{(\rho)}(\bm{x}_{t_{n}}|\bm{x}_{0},\bm{x}_{t_{n+1}},\bm{x}_{T})$ is defined in Eqn. (11) as
@@
C.2 The Insignificance of Loss Weighting in Training
The insignificance of the weighting mismatch in Proposition 3.2 can be interpreted from two aspects. On the one hand, $\mathcal{L}$ consists of independent terms concerning individual timesteps (as long as the network’s parameters are not shared across different $t$), ensuring that the global minimum remains the same as minimizing the loss at each timestep, regardless of the weighting. On the other hand, $\mathcal{L}$ under different weightings are mutually bounded by $\frac{\min_{t}w_{t}}{\max_{t}\gamma_{t}}\mathcal{L}_{\gamma}(\theta)\leq\mathcal{L}_{w}(\theta)\leq\frac{\max_{t}w_{t}}{\min_{t}\gamma_{t}}\mathcal{L}_{\gamma}(\theta)$. Besides, it is widely acknowledged that in diffusion models, the weighting corresponding to variational inference may yield superior likelihood but suboptimal sample quality (Ho et al., 2020; Song et al., 2021c), which is not preferred in practice.
C.3 Special Cases and Relationship with Prior Works
Connection to Flow Matching
When the noise schedule $\alpha_{t}=1,T=1$ and $\sigma_{t}=\sqrt{\beta t}$, the forward process becomes $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{1})=\mathcal{N}(t\bm{x}_{T}+(1-t)\bm{x}_{0},\beta t(1-t))$ which is a Brownian bridge. When $\beta\rightarrow 0$, there will be no intermediate noise and the forward process is similar to flow matching (Lipman et al., 2022). In this limit, the DBIM ($\eta=1$) updating rule from time $t$ to time $s<t$ will become $\bm{x}_{s}=s\bm{x}_{T}+(1-s)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\frac{s}{t}\bm{x}_{t}+(1-\frac{s}{t})\bm{x}_{\theta}(\bm{x}_{t},t)=\bm{x}_{t}-(t-s)\bm{v}_{\theta}(\bm{x}_{t},t)$. Here we define $\bm{v}_{\theta}(\bm{x}_{t},t)\coloneqq\frac{\bm{x}_{t}-\bm{x}_{\theta}(\bm{x}_{t},t)}{t}$ as the velocity function of the probability flow (i.e., the drift of the ODE) in flow matching methods. Therefore, in the flow matching case, DBIM is a simple Euler step of the flow.
Connection to DDIM
In the regions where $t$ is small and $\frac{\text{SNR}_{T}}{\text{SNR}_{t}}$ is close to 0, we have $a_{t}\approx 0,b_{t}\approx\alpha_{t},c_{t}\approx\sigma_{t}$. Therefore, the forward process of DDBM in this case is approximately $q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{T})=\mathcal{N}(\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$, which is the forward process of the corresponding diffusion model. Moreover, in this case, the DBIM ($\eta=0$) step is approximately
$\bm{x}_{s}\approx\frac{\sigma_{s}}{\sigma_{t}}\bm{x}_{t}+\sigma_{s}\left(\frac{\alpha_{s}}{\sigma_{s}}-\frac{\alpha_{t}}{\sigma_{t}}\right)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ (52)
which is exactly DDIM (Song et al., 2021a), except that the data prediction network $\bm{x}_{\theta}$ is dependent on $\bm{x}_{T}$. This indicates that when $t$ is small so that the component of $x_{T}$ in $x_{t}$ is negligible, DBIM recovers DDIM.
Connection to Posterior Sampling
The previous work I2SB (Liu et al., 2023) also employs diffusion bridges with discrete timesteps, though their noise schedule is restricted to the variance exploding (VE) type with $f(t)=0$ in the forward process. For generation, they adopt a similar approach to DDPM (Ho et al., 2020) by iterative sampling from the posterior distribution $p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n})$, which is a parameterized and shortened diffusion bridge between the endpoints $\hat{\bm{x}}_{0}=\bm{x}_{\theta}(\bm{x}_{n},t_{n},\bm{x}_{N})$ and $\bm{x}_{n}$. Since the posterior distribution is not conditioned on $\bm{x}_{T}$ (except through the parameterized network), the corresponding forward diffusion bridge is Markovian. Thus, the posterior sampling in I2SB is a special case of DBIM by setting $\eta=0$ and $f(t)=0$.
C.4 Perspective of Exponential Integrators
Exponential integrators (Calvo & Palencia, 2006; Hochbruck et al., 2009) are widely adopted in recent works concerning fast sampling of diffusion models (Zhang & Chen, 2022; Zheng et al., 2023a; Gonzalez et al., 2023). Suppose we have an ODE
$\mathrm{d}\bm{x}_{t}=[a(t)\bm{x}_{t}+b(t)\bm{F}_{\theta}(\bm{x}_{t},t)]\mathrm{d}t$ (53)
where $\bm{F}_{\theta}$ is the parameterized prediction function that we want to approximate with Taylor expansion. The usual way of representing its analytic solution $\bm{x}_{t}$ at time $t$ with respect to an initial condition $\bm{x}_{s}$ at time $s$ is
$\bm{x}_{t}=\bm{x}_{s}+\int_{s}^{t}[a(\tau)\bm{x}_{\tau}+b(\tau)\bm{F}_{\theta}(\bm{x}_{\tau},\tau)]\mathrm{d}\tau$ (54)
By approximating the involved integrals in Eqn. (54), we can obtain direct discretizations of Eqn. (53) such as Euler’s method. The key insight of exponential integrators is that, it is often better to utilize the “semi-linear” structure of Eqn. (53) and analytically cancel the linear term $a(t)\bm{x}_{t}$. This way, we can obtain solutions that only involve integrals of $\bm{F}_{\theta}$ and result in lower discretization errors. Specifically, by the “variation-of-constants” formula, the exact solution of Eqn. (53) can be alternatively given by
$\bm{x}_{t}=e^{\int_{s}^{t}a(\tau)\mathrm{d}\tau}\bm{x}_{s}+\int_{s}^{t}e^{\int_{\tau}^{t}a(r)\mathrm{d}r}b(\tau)\bm{F}_{\theta}(\bm{x}_{\tau},\tau)\mathrm{d}\tau$ (55)
We can apply this transformation to the PF-ODE in DDBMs. By collecting the linear terms w.r.t. $\bm{x}_{t}$, Eqn. (8) can be rewritten as (already derived in Appendix B.3)
$\mathrm{d}\bm{x}_{t}=\left[\left(f(t)+\frac{g^{2}(t)}{\sigma_{t}^{2}}-\frac{g^{2}(t)}{2c_{t}^{2}}\right)\bm{x}_{t}+g^{2}(t)\frac{a_{t}}{2c_{t}^{2}}\bm{x}_{T}-g^{2}(t)\frac{b_{t}}{2c_{t}^{2}}\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})\right]\mathrm{d}t$ (56)
By corresponding it to Eqn. (53), we have
$a(t)=f(t)+\frac{g^{2}(t)}{\sigma_{t}^{2}}-\frac{g^{2}(t)}{2c_{t}^{2}},\quad b_{1}(t)=g^{2}(t)\frac{a_{t}}{2c_{t}^{2}},\quad b_{2}(t)=-g^{2}(t)\frac{b_{t}}{2c_{t}^{2}}$ (57)
From Eqn. (43), Eqn. (46) and Eqn. (47), we know
$a(t)=\frac{\mathrm{d}\log c_{t}}{\mathrm{d}t},\quad b_{1}(t)=a_{t}\frac{\mathrm{d}\log(a_{t}/c_{t})}{\mathrm{d}t},\quad b_{2}(t)=b_{t}\frac{\mathrm{d}\log(b_{t}/c_{t})}{\mathrm{d}t}$ (58)
Note that these relations are known in advance when converting from our ODE to the PF-ODE. Otherwise, finding them in this inverse conversion will be challenging. The integrals in Eqn. (55) can then be calculated as $\int_{s}^{t}a(\tau)\mathrm{d}\tau=\log c_{t}-\log c_{s}$. Thus
$e^{\int_{s}^{t}a(\tau)\mathrm{d}\tau}=\frac{c_{t}}{c_{s}},\quad e^{\int_{\tau}^{t}a(r)\mathrm{d}r}=\frac{c_{t}}{c_{\tau}}$ (59)
Therefore, the exact solution in Eqn. (55) becomes
$\bm{x}_{t}$ $=\frac{c_{t}}{c_{s}}\bm{x}_{s}+c_{t}\int_{s}^{t}\frac{a_{\tau}}{c_{\tau}}\bm{x}_{T}\mathrm{d}\log\left(\frac{a_{\tau}}{c_{\tau}}\right)+c_{t}\int_{s}^{t}\frac{b_{\tau}}{c_{\tau}}\bm{x}_{\theta}(\bm{x}_{\tau},\tau,\bm{x}_{T})\mathrm{d}\log\left(\frac{b_{\tau}}{c_{\tau}}\right)$ (60)
$=\frac{c_{t}}{c_{s}}\bm{x}_{s}+\left(a_{t}-\frac{c_{t}}{c_{s}}a_{s}\right)\bm{x}_{T}+c_{t}\int_{s}^{t}\frac{b_{\tau}}{c_{\tau}}\bm{x}_{\theta}(\bm{x}_{\tau},\tau,\bm{x}_{T})\mathrm{d}\log\left(\frac{b_{\tau}}{c_{\tau}}\right)$
which is the same as Eqn. (17) after exchanging $s$ and $t$ and changing the time variable in the integral to $\lambda_{t}=\log\left(\frac{b_{t}}{c_{t}}\right)$.
Lastly, we emphasize the advantage of DBIM over employing exponential integrators. First, deriving our ODE via exponential integrators requires the PF-ODE as preliminary. However, the PF-ODE alone cannot handle the singularity at the start point and presents theoretical challenges. Moreover, the conversion process from the PF-ODE to our ODE is intricate, while DBIM retains the overall simplicity. Additionally, DBIM supports varying levels of stochasticity during sampling, unlike the deterministic nature of ODE-based methods. This stochasticity can mitigate sampling errors via the Langevin mechanism (Song et al., 2021c), potentially enhancing the generation quality.
Appendix D Derivation of Our High-Order Numerical Solvers
High-order solvers of Eqn. (17) can be developed by approximating $\bm{x}_{\theta}$ in the integral with Taylor expansion. Specifically, as a function of $\lambda$, we have $\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\approx\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+\sum_{k=1}^{n}\frac{(\lambda-\lambda_{t})^{k}}{k!}\bm{x}^{(k)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$, where $\bm{x}^{(k)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\frac{\mathrm{d}^{k}\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})}{\mathrm{d}\lambda^{k}}\Big{|}_{\lambda=\lambda_{t}}$ is the $k$-th order derivative w.r.t. $\lambda$, which can be estimated with finite difference of past network outputs.
2nd-Order Case
With the Taylor expansion $\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\approx\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+(\lambda-\lambda_{t})\bm{x}^{(1)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$, we have
$\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\mathrm{d}\lambda$ $\approx\left(\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\mathrm{d}\lambda\right)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+\left(\int_{\lambda_{t}}^{\lambda_{s}}(\lambda-\lambda_{t})e^{\lambda}\mathrm{d}\lambda\right)\bm{x}^{(1)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ (61)
$\approx e^{\lambda_{s}}\left[(1-e^{-h})\hat{\bm{x}}_{t}+(h-1+e^{-h})\hat{\bm{x}}_{t}^{(1)}\right]$
where we use $\hat{\bm{x}}_{t}$ to denote the network output at time $t$, and $h=\lambda_{s}-\lambda_{t}>0$. Suppose we have used a previous timestep $u$ ($s<t<u$), the first-order derivative can be estimated by
$\hat{\bm{x}}_{t}^{(1)}\approx\frac{\hat{\bm{x}}_{t}-\hat{\bm{x}}_{u}}{h_{1}},\quad h_{1}=\lambda_{t}-\lambda_{u}$ (62)
3rd-Order Case
With the Taylor expansion $\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\approx\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+(\lambda-\lambda_{t})\bm{x}^{(1)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+\frac{(\lambda-\lambda_{t})^{2}}{2}\bm{x}^{(2)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$, we have
$\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\bm{x}_{\theta}(\bm{x}_{t_{\lambda}},t_{\lambda},\bm{x}_{T})\mathrm{d}\lambda$ (63)
$\approx$ $\left(\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\mathrm{d}\lambda\right)\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})+\left(\int_{\lambda_{t}}^{\lambda_{s}}(\lambda-\lambda_{t})e^{\lambda}\mathrm{d}\lambda\right)\bm{x}^{(1)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$
$+\left(\int_{\lambda_{t}}^{\lambda_{s}}\frac{(\lambda-\lambda_{t})^{2}}{2}e^{\lambda}\mathrm{d}\lambda\right)\bm{x}^{(2)}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$
$\approx$ $e^{\lambda_{s}}\left[(1-e^{-h})\hat{\bm{x}}_{t}+(h-1+e^{-h})\hat{\bm{x}}_{t}^{(1)}+\left(\frac{h^{2}}{2}-h+1-e^{-h}\right)\hat{\bm{x}}_{t}^{(2)}\right]$
Similarly, suppose we have two previous timesteps $u_{1},u_{2}$ ($s<t<u_{1}<u_{2}$), and denote $h_{1}\coloneqq\lambda_{t}-\lambda_{u_{1}},h_{2}\coloneqq\lambda_{u_{1}}-\lambda_{u_{2}}$, the first-order and second-order derivatives can be estimated by
$\hat{\bm{x}}_{t}^{(1)}\approx\frac{\frac{\hat{\bm{x}}_{t}-\hat{\bm{x}}_{u_{1}}}{h_{1}}(2h_{1}+h_{2})-\frac{\hat{\bm{x}}_{u_{1}}-\hat{\bm{x}}_{u_{2}}}{h_{2}}h_{1}}{h_{1}+h_{2}},\quad\hat{\bm{x}}_{t}^{(2)}\approx 2\frac{\frac{\hat{\bm{x}}_{t}-\hat{\bm{x}}_{u_{1}}}{h_{1}}-\frac{\hat{\bm{x}}_{u_{1}}-\hat{\bm{x}}_{u_{2}}}{h_{2}}}{h_{1}+h_{2}}$ (64)
The high-order samplers for DDBMs also theoretically guarantee the order of convergence, similar to those for diffusion models (Zhang & Chen, 2022; Lu et al., 2022b; Zheng et al., 2023a). We omit the proofs here as they deviate from our main contributions.
Appendix E Algorithm
Algorithm 1 DBIM (high-order)
1:condition $\bm{x}_{T}$, timesteps $0\leq t_{0}<t_{1}<\dots<t_{N-1}<t_{N}=T$, data prediction model $\bm{x}_{\theta}$, booting noise $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, noise schedule $a_{t},b_{t},c_{t}$, $\lambda_{t}=\log(b_{t}/c_{t})$, order $o$ (2 or 3).
2:$\hat{\bm{x}}_{T}\leftarrow\bm{x}_{\theta}(\bm{x}_{T},T,\bm{x}_{T})$
3:$\bm{x}_{t_{N-1}}\leftarrow a_{t}\bm{x}_{T}+b_{t}\hat{\bm{x}}_{T}+c_{t}\bm{\epsilon}$
4:for $i\leftarrow N-1$ to $1$ do
5:     $s,t\leftarrow t_{i-1},t_{i}$; $h\leftarrow\lambda_{s}-\lambda_{t}$
6:     $\hat{\bm{x}}_{t}\leftarrow\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$
7:     if $o=2$ or $i=N-1$ then
8:         $u\leftarrow t_{i+1}$; $h_{1}\leftarrow\lambda_{t}-\lambda_{u}$
9:         Estimate $\hat{\bm{x}}_{t}^{(1)}$ with Eqn. (62)
10:         $\hat{\bm{I}}\leftarrow e^{\lambda_{s}}\left[(1-e^{-h})\hat{\bm{x}}_{t}+(h-1+e^{-h})\hat{\bm{x}}_{t}^{(1)}\right]$
11:     else
12:         $u_{1},u_{2}\leftarrow t_{i+1},t_{i+2}$; $h_{1}\leftarrow\lambda_{t}-\lambda_{u_{1}}$; $h_{2}\leftarrow\lambda_{u_{1}}-\lambda_{u_{2}}$
13:         Estimate $\hat{\bm{x}}_{t}^{(1)},\hat{\bm{x}}_{t}^{(2)}$ with Eqn. (64)
14:         $\hat{\bm{I}}\leftarrow e^{\lambda_{s}}\left[(1-e^{-h})\hat{\bm{x}}_{t}+(h-1+e^{-h})\hat{\bm{x}}_{t}^{(1)}+\left(\frac{h^{2}}{2}-h+1-e^{-h}\right)\hat{\bm{x}}_{t}^{(2)}\right]$
15:     end if
16:     $\bm{x}_{s}\leftarrow\frac{c_{s}}{c_{t}}\bm{x}_{t}+\left(a_{s}-\frac{c_{s}}{c_{t}}a_{t}\right)\bm{x}_{T}+c_{s}\hat{\bm{I}}$
17:end for
18:return $\bm{x}_{t_{0}}$
Appendix F Experiment Details
F.1 Model Details